# Optimal Control of Reserve Asset Portfolios for Pegged Digital Currencies

### v1

@@ 3.2 Event-Triggered Optimization via Pontryagin’s Maximum Principle
$\quad+p_{\Delta P}\left(\eta\frac{[\hat{Q}-R_{\mathrm{liq}}]_{+}}{\hat{S}_{\mathrm{out}}}-\gamma\delta\right),$ (28)
the costates satisfy (almost everywhere at points where $R_{\mathrm{liq}}\neq\hat{Q}$)
$\dot{p}_{\mathrm{liq}}$ $=\rho p_{\mathrm{liq}}-\frac{\partial H}{\partial R_{\mathrm{liq}}}=\rho p_{\mathrm{liq}}-\left(p_{\mathrm{liq}}r_{\mathrm{cash}}-p_{\Delta P}\eta\frac{\mathbf{1}_{\{\hat{Q}>R_{\mathrm{liq}}\}}}{\hat{S}_{\mathrm{out}}}\right),$ (29)
$\dot{p}_{\mathrm{bill}}$ $=\rho p_{\mathrm{bill}}-\frac{\partial H}{\partial R_{\mathrm{bill}}}=\rho p_{\mathrm{bill}}-(-r_{\mathrm{bill}}+p_{\mathrm{bill}}r_{\mathrm{bill}})=(\rho-r_{\mathrm{bill}})p_{\mathrm{bill}}+r_{\mathrm{bill}},$ (30)
$\dot{p}_{\Delta P}$ $=\rho p_{\Delta P}-\frac{\partial H}{\partial\Delta P}=\rho p_{\Delta P}-2c_{\mathrm{peg}}\Delta P,$ (31)
where $\mathbf{1}_{\{\hat{Q}>R_{\mathrm{liq}}\}}$ is the indicator function that equals 1 when redemptions exceed liquid reserves and 0 otherwise.
@@
where $\mathrm{shrink}(z,\lambda)=\mathrm{sign}(z)\,\max\{|z|-\lambda,\,0\}$ denotes the soft-thresholding operator. This implies that the optimal portfolio reallocation control $\omega$ at time $t$ follows a threshold policy: if the absolute marginal benefit of reallocation does not exceed the transaction cost $(|S_{\omega}(t)|\leq\lambda_{\omega})$, no reallocation within the reserve asset portfolio occurs $(\omega^{*}(t)=0)$. If the marginal benefit exceeds the transaction cost $(|S_{\omega}(t)|>\lambda_{\omega})$, reallocation is performed at the maximal admissible rate $(\omega^{*}(t)\in\{-\omega_{\max},\,\omega_{\max}\})$.
@@ 3.3.1 A Discretized Optimal Control Policy
$\int_{\tau_{j}}^{\tau_{j+1}}\left(\lambda_{\omega}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\omega_{j}^{2}-(p_{\mathrm{bill}}(t)-p_{\mathrm{liq}}(t))\omega_{j}\right)dt.$ (38)
@@
$J_{j}(\omega_{j})=\int_{\tau_{j}}^{\tau_{j+1}}\left(\lambda_{\omega}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\omega_{j}^{2}-(p_{\mathrm{bill}}(t)-p_{\mathrm{liq}}(t))\omega_{j}\right)dt$ (42)
@@
$J_{j}(\omega_{j})=\lambda_{\omega}\Delta_{j}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\Delta_{j}\omega_{j}^{2}-S_{j}\omega_{j},$ (43)
@@
$0\in\lambda_{\omega}\Delta_{j}\cdot\partial|\omega_{j}|+\rho_{\omega}\Delta_{j}\omega_{j}-S_{j}.$ (45)
@@
$0=\lambda_{\omega}\Delta_{j}+\rho_{\omega}\Delta_{j}\omega_{j}-S_{j}\implies\omega_{j}=\frac{S_{j}-\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}.$ (46)
This is valid only if $\omega_{j}>0$, i.e., when $S_{j}>\lambda_{\omega}\Delta_{j}$.
@@
$0=-\lambda_{\omega}\Delta_{j}+\rho_{\omega}\Delta_{j}\omega_{j}-S_{j}\implies\omega_{j}=\frac{S_{j}+\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}.$ (47)
This is valid only if $\omega_{j}<0$, i.e., when $S_{j}<-\lambda_{\omega}\Delta_{j}$.
@@
$0\in\lambda_{\omega}\Delta_{j}[-1,1]+0-S_{j}=[-\lambda_{\omega}\Delta_{j}-S_{j},\lambda_{\omega}\Delta_{j}-S_{j}].$ (48)
@@
$\omega_{j}^{\text{unc}}=\begin{cases}\frac{S_{j}-\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}&\text{if }S_{j}>\lambda_{\omega}\Delta_{j}\\ 0&\text{if }|S_{j}|\leq\lambda_{\omega}\Delta_{j}\\ \frac{S_{j}+\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}&\text{if }S_{j}<-\lambda_{\omega}\Delta_{j}\end{cases}$ (49)
@@ 3.3.2 Monotone Stress-Response Behavior
    $\omega_{j}^{*}(\hat{Q}_{\alpha})=\begin{cases}0&\text{if }\alpha\leq\alpha^{*}\\ \in\{-\omega_{\max},0,+\omega_{\max}\}&\text{if }\alpha>\alpha^{*}\end{cases}$ (52)
@@
For the peg deviation, since $\cdot$$_{+}$ is monotone increasing and $\hat{Q}^{(1)}-R_{\mathrm{liq}}^{(1)}\geq\hat{Q}^{(2)}-R_{\mathrm{liq}}^{(2)}$:
$\dot{\Delta P}^{(1)}-\dot{\Delta P}^{(2)}=\eta\frac{[\hat{Q}^{(1)}-R_{\mathrm{liq}}^{(1)}]_{+}-[\hat{Q}^{(2)}-R_{\mathrm{liq}}^{(2)}]_{+}}{\hat{S}_{\mathrm{out}}}\geq 0,$ (54)
yielding $\Delta P^{(1)}(t)\geq\Delta P^{(2)}(t)$.
@@
$p_{\Delta P}(t)=e^{\rho(t-\tau_{j+1})}p_{\Delta P}(\tau_{j+1})-2c_{\mathrm{peg}}\int_{t}^{\tau_{j+1}}e^{\rho(t-s)}\Delta P(s)\,ds.$ (55)
Since $\Delta P^{(1)}\geq\Delta P^{(2)}$, we have $p_{\Delta P}^{(1)}(t)\leq p_{\Delta P}^{(2)}(t)$.
@@
$\dot{p}_{\mathrm{liq}}=(\rho-r_{\mathrm{cash}})p_{\mathrm{liq}}+\eta\frac{\mathbf{1}_{\{\hat{Q}>R_{\mathrm{liq}}\}}}{\hat{S}_{\mathrm{out}}}p_{\Delta P}$ (56)
form a linear system with monotone coefficient (the indicator function) and monotone forcing (since $p_{\Delta P}^{(1)}\leq p_{\Delta P}^{(2)}$).Hence,
@@ 4.1 Experimental Design
We conduct numerical experiments over 92 calendar days (August 1 - October 31, 2025) with three daily settlement windows at 10:00, 13:00, and 16:00 ET, aligning with major settlement times. The performance of various control strategies is evaluated using three scenarios that capture characteristic market stress patterns empirically observed in financial markets. The construction of these scenarios enables precise testing of our central hypothesis that the developed control algorithm fulfills three essential requirements: first, the exploitation of lucrative investment opportunities in Treasury Bills during calm market phases; second, the maintenance of sufficient liquidity reserves in volatile phases with potentially endogenous bank-run dynamics; and third, the early detection of emerging volatility phases, such that proactive liquidity procurement is initiated at early signs of a demand shock. To systematically investigate these properties, we define the following market scenarios:
@@
    Single shock event: The system operates under baseline conditions for days 0-30, experiences elevated redemption pressure during days 30-45 with baseline intensity $\lambda_{0,R}$ multiplied by a factor $\chi\sim\text{Uniform}[2,4]$, then undergoes exponential mean-reversion with a 10-day time constant through day 92. This stress test examines the controller’s ability to respond decisively to sustained redemption pressure and subsequently restore normalized operations.
@@
Peg stability is measured by the failure probability, defined as the relative frequency of complete depeg events ($\Delta P=1$),. Responsiveness is measured by the time lag between a shock onset and the controller’s increase in cash-buying speed.
In addition to the optimal control described in Subsection 3.3, we evaluate the performance of two reference strategies:
@@ 4.2 Parameter Estimation
The interest rates used in our numerical computations are taken from current yield curves and Federal Reserve projections, expressed on a per-settlement-window basis (8 hours):
  • 90-day Treasury rate: $r_{\mathrm{bill}}=0.054/(3\times 365)\approx 4.93\times 10^{-5}$ per window (5.4% annualized yield as of August 2025)
  • Discount rate: $\rho=0.08/(3\times 365)\approx 7.31\times 10^{-5}$ per window (8% annualized cost of capital reflecting regulatory risk premium)
  • Discount rate: $\rho=0.08\times 3/8760\approx 2.74\times 10^{-5}$ per window (8% annualized cost of capital reflecting regulatory risk premium)
@@
  • Maximum reallocation rate: $\omega_{\max}=0.1\times S_{\mathrm{out}}$ per window (10% of supply)
@@
To simplify the interpretation of numerical results, we assume zero mint or redemption fee ($\delta=0$). For modelling the Hawkes processes, we choose parameters to mimic the dynamics observed in major stablecoins (USDC, USDT) during 2024-2025. For baseline intensities, we set $\lambda_{0,R}=100$ events/hour for redemptions and $\lambda_{0,M}=80$ events/hour for mints, corresponding to daily baseline volumes of 2,400 and 1,920 events respectively, which align with typical median volumes for these assets.
@@
To capture the clustering behavior characteristic of stablecoin transactions, we parameterize the self-excitation dynamics with decay rates of $\theta_{R}=2.0$ hour-1 and $\theta_{M}=1.5$ hour-1, yielding half-lives of approximately 20-30 minutes that reflect how quickly bursts of activity dissipate. We set the corresponding jump sizes to $\kappa_{R}=0.8$ and $\kappa_{M}=0.6$, which ensure subcritical branching ratios ($\kappa_{R}/\theta_{R}=0.4<1$ and $\kappa_{M}/\theta_{M}=0.4<1$) necessary for process stability. Finally, transaction sizes follow exponential distributions with mean redemption size $\bar{z}_{R}=\$250,000$ and mean mint size $\bar{z}_{M}=\$300,000$, values representative of institutional trading volumes in these markets.
The cost function parameters are derived from economic fundamentals and regulatory objectives. For peg stability, we assume that deviations from the dollar peg create arbitrage opportunities proportional to $(\Delta P)^{2}$. To maintain deviations below 10 basis points under normal conditions, we calibrate the peg stability cost as:
$c_{\mathrm{peg}}=\frac{\text{Daily carry value}}{\text{Tolerance}^{2}}=\frac{r_{\mathrm{bill}}\times S_{\mathrm{out}}/365}{(0.001)^{2}}\approx 1.5\times 10^{8}$ (59)
This ensures that the controller’s prioritization of peg maintenance is commensurate with the economic value at stake. The fee penalty parameter balances user experience with revenue requirements, calibrated to target fees around 10 basis points:
$c_{\mathrm{fee}}=\frac{\text{Median daily redemption value}}{(\text{Target fee})^{2}}=\frac{\lambda_{0,R}\times\bar{z}_{R}\times 24}{(0.01)^{2}}\approx 6\times 10^{8}$ (60)
Trading costs reflect Treasury market microstructure, with a linear component $\lambda_{\omega}=0.0001\times S_{\mathrm{out}}$ representing the 1 basis point bid-ask spread and a quadratic component $\rho_{\omega}=0.00001\times S_{\mathrm{out}}$ capturing price impact for large trades. The peg dynamics are parameterized based on historical stablecoin de-pegging events: we set the supply-peg sensitivity $\eta=10$, calibrated so that a 10% supply shortfall creates a 1% peg deviation; the fee effectiveness parameter $\gamma=5$, whereby a 50 basis point fee counters a 10 basis point deviation; and the peg feedback strength $\zeta=100$, such that a 1% deviation doubles redemption intensity.
@@
For initial conditions, we consider a system at scale with outstanding supply $S_{\mathrm{out}}(0)=\$10$ billion and zero deviation from the peg ($\Delta P(0)=0$). Transaction intensities begin at baseline levels: $\lambda_{R}(0)=\lambda_{0,R}$ and $\lambda_{M}(0)=\lambda_{0,M}$. The reserve allocation strategies differ between controllers: both the optimal controller and the maximum yield controller begin with $R_{\mathrm{liq}}(0)=\$1$ billion (10%) in liquid assets and $R_{\mathrm{bill}}(0)=\$9$ billion (90%) in Treasury bills, while the maximum liquidity controller starts with all reserves (\$10 billion) allocated to liquid cash.

### v2

@@ 3.2 Optimal Control Characterization via Pontryagin’s Maximum Principle
$\quad+p_{\Delta P}\!\left(\eta\!\left(\mathbf{1}_{\{\hat{S}_{\mathrm{out}}>0\}}\!\left[\frac{\tau_{\mathrm{set}}[\hat{Q}]_{+}-R_{\mathrm{liq}}}{\hat{S}_{\mathrm{out}}}\right]_{0}^{1}\!-\Delta P\right)-\gamma\,\delta\right),$ (36)
the costates satisfy (almost everywhere at points where $R_{\mathrm{liq}}\neq\hat{S}_{\mathrm{out}}$)
$\dot{p}_{\mathrm{liq}}$ $=\rho p_{\mathrm{liq}}-\frac{\partial H}{\partial R_{\mathrm{liq}}}=(\rho-r_{\mathrm{cash}})p_{\mathrm{liq}}+p_{\Delta P}\,\eta\,\frac{\mathbf{1}_{\{\tau_{\mathrm{set}}[\hat{Q}(t)]_{+}>R_{\mathrm{liq}}\}}}{\hat{S}_{\mathrm{out}}(t)},$ (37)
$\dot{p}_{\mathrm{bill}}$ $=\rho p_{\mathrm{bill}}-\frac{\partial H}{\partial R_{\mathrm{bill}}}=(\rho-r_{\mathrm{bill}})p_{\mathrm{bill}}+r_{\mathrm{bill}},$ (38)
$\dot{p}_{\Delta P}$ $=\rho p_{\Delta P}-\frac{\partial H}{\partial\Delta P}=(\rho+\eta)\,p_{\Delta P}-2c_{\mathrm{peg}}\,\Delta P,$ (39)
where $\mathbf{1}_{\{\tau_{\mathrm{set}}[\hat{Q}(t)]_{+}>R_{\mathrm{liq}}\}}$ indicates a forecast cash shortfall over the settlement horizon.
@@
where $\mathrm{shrink}(z,\lambda)=\mathrm{sign}(z)\,\max\{|z|-\lambda,\,0\}$ denotes the soft-thresholding operator. This implies that the optimal portfolio reallocation control $\omega$ at time $t$ follows a threshold policy: if the absolute marginal benefit of reallocation does not exceed the transaction cost $(|S_{\omega}(t)|\leq\lambda_{\omega})$, no reallocation within the reserve asset portfolio occurs $(\omega^{*}(t)=0)$. If the marginal benefit exceeds the transaction-cost threshold $(|S_{\omega}(t)|>\lambda_{\omega})$, reallocation occurs with magnitude increasing in $|S_{\omega}(t)|-\lambda_{\omega}$ and capped at $\omega_{\max}$ by the projection, i.e. $0<|\omega^{*}(t)|\leq\omega_{\max}$ with $\mathrm{sign}(\omega^{*}(t))=-\mathrm{sign}(S_{\omega}(t))$.
@@ 3.3.1 A Discretized Optimal Control Policy
$\int_{\tau_{j}}^{\tau_{j+1}}\left(\lambda_{\omega}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\omega_{j}^{2}+(p_{\mathrm{bill}}(t)-p_{\mathrm{liq}}(t))\omega_{j}\right)dt.$ (46)
@@
$J_{j}(\omega_{j})=\int_{\tau_{j}}^{\tau_{j+1}}\left(\lambda_{\omega}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\omega_{j}^{2}+(p_{\mathrm{bill}}(t)-p_{\mathrm{liq}}(t))\omega_{j}\right)dt$ (50)
@@
$J_{j}(\omega_{j})=\lambda_{\omega}\Delta_{j}|\omega_{j}|+\tfrac{1}{2}\rho_{\omega}\Delta_{j}\omega_{j}^{2}+S_{j}\omega_{j},$ (51)
@@
$0\in\lambda_{\omega}\Delta_{j}\cdot\partial|\omega_{j}|+\rho_{\omega}\Delta_{j}\omega_{j}+S_{j}.$ (53)
@@
$0=\lambda_{\omega}\Delta_{j}+\rho_{\omega}\Delta_{j}\omega_{j}+S_{j}\implies\omega_{j}=\frac{-S_{j}-\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}.$ (54)
This is valid only if $\omega_{j}>0$, i.e., when $S_{j}<-\lambda_{\omega}\Delta_{j}$.
@@
$0=-\lambda_{\omega}\Delta_{j}+\rho_{\omega}\Delta_{j}\omega_{j}+S_{j}\implies\omega_{j}=\frac{-S_{j}+\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}.$ (55)
This is valid only if $\omega_{j}<0$, i.e., when $S_{j}>\lambda_{\omega}\Delta_{j}$.
@@
$0\in\lambda_{\omega}\Delta_{j}[-1,1]+0+S_{j}=[-\lambda_{\omega}\Delta_{j}+S_{j},\lambda_{\omega}\Delta_{j}+S_{j}].$ (56)
@@
$\omega_{j}^{\text{unc}}=\begin{cases}\frac{-S_{j}-\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}&\text{if }S_{j}<-\lambda_{\omega}\Delta_{j}\\[4.0pt] 0&\text{if }|S_{j}|\leq\lambda_{\omega}\Delta_{j}\\[4.0pt] \frac{-S_{j}+\lambda_{\omega}\Delta_{j}}{\rho_{\omega}\Delta_{j}}&\text{if }S_{j}>\lambda_{\omega}\Delta_{j}\end{cases}$ (57)
@@ 3.3.2 Monotone Stress-Response Behavior
    $\omega_{j}^{*}(\hat{Q}_{\alpha})=\begin{cases}0&\text{if }\alpha\leq\alpha^{*}\\ \neq 0&\text{if }\alpha>\alpha^{*}\end{cases}$ (60)
@@
For the peg deviation, the fair-value deviations are
$\Delta P_{\mathrm{fair}}^{(i)}(t):=\mathbf{1}_{\{\hat{S}_{\mathrm{out}}(t)>0\}}\left[\frac{\tau_{\mathrm{set}}[\hat{Q}^{(i)}(t)]_{+}-R_{\mathrm{liq}}^{(i)}(t)}{\hat{S}_{\mathrm{out}}(t)}\right]_{0}^{1},\qquad i\in\{1,2\}.$
Since $\hat{Q}^{(1)}(t)\geq\hat{Q}^{(2)}(t)$ and $R_{\mathrm{liq}}^{(1)}(t)\leq R_{\mathrm{liq}}^{(2)}(t)$, the numerator $\tau_{\mathrm{set}}[\hat{Q}^{(i)}]_{+}-R_{\mathrm{liq}}^{(i)}$ is weakly larger for $i=1$ than for $i=2$, and the $\cdot]_{0}^{1}$ clipping preserves this ordering, giving $\Delta P_{\mathrm{fair}}^{(1)}(t)\geq\Delta P_{\mathrm{fair}}^{(2)}(t)$. Under identical fee controls $\delta_{j}$, the peg dynamics imply that the difference $D(t):=\Delta P^{(1)}(t)-\Delta P^{(2)}(t)$ satisfies
$\dot{D}(t)=\eta\big(\Delta P_{\mathrm{fair}}^{(1)}(t)-\Delta P_{\mathrm{fair}}^{(2)}(t)\big)-\eta D(t).$ (62)
With $D(\tau_{j})=0$ and nonnegative forcing, comparison yields $D(t)\geq 0$ for all $t\in[\tau_{j},\tau_{j+1})$, i.e. $\Delta P^{(1)}(t)\geq\Delta P^{(2)}(t)$.
@@
$p_{\Delta P}(t)=e^{(\rho+\eta)(t-\tau_{j+1})}p_{\Delta P}(\tau_{j+1})+2c_{\mathrm{peg}}\int_{t}^{\tau_{j+1}}e^{(\rho+\eta)(t-s)}\Delta P(s)\,ds.$ (63)
Since $\Delta P^{(1)}\geq\Delta P^{(2)}$, we have $p_{\Delta P}^{(1)}(t)\geq p_{\Delta P}^{(2)}(t)$.
@@
$\dot{p}_{\mathrm{liq}}=(\rho-r_{\mathrm{cash}})p_{\mathrm{liq}}+\eta\,\frac{\mathbf{1}_{\{\tau_{\mathrm{set}}[\hat{Q}(t)]_{+}>R_{\mathrm{liq}}\}}}{\hat{S}_{\mathrm{out}}}\,p_{\Delta P}$ (64)
form a linear system with monotone coefficient (the indicator function) and monotone forcing (since $p_{\Delta P}^{(1)}\geq p_{\Delta P}^{(2)}$ and $p_{\Delta P}\geq 0$ for $\Delta P\geq 0$). Hence,
@@ 4.1 Experimental Design
We conduct numerical experiments over 92 calendar days (August 1 - October 31, 2025) with three equally spaced settlement windows per day (so $\Delta_{j}=8$ hours), aligning with major settlement times. The performance of various control strategies is evaluated using three scenarios that capture characteristic market stress patterns empirically observed in financial markets. The construction of these scenarios enables precise testing of our central hypothesis that the developed control algorithm fulfills three essential requirements: first, the exploitation of lucrative investment opportunities in Treasury Bills during calm market phases; second, the maintenance of sufficient liquidity reserves in volatile phases with potentially endogenous bank-run dynamics; and third, the early detection of emerging volatility phases, such that proactive liquidity procurement is initiated at early signs of a demand shock. To systematically investigate these properties, we define the following market scenarios:
@@
    Single shock event: The system operates under baseline conditions for days 0-30, experiences elevated redemption pressure during days 30-45 with baseline intensity $\lambda_{0,R}$ multiplied by a factor $\chi\sim\text{Uniform}[2,4]$, then undergoes exponential mean-reversion with a 4-day time constant through day 92. This stress test examines the controller’s ability to respond decisively to sustained redemption pressure and subsequently restore normalized operations.
@@
Peg stability is measured by the failure probability, defined as the relative frequency of severe depeg events ($\Delta P\geq 0.85$). Responsiveness is measured by the time lag between a shock onset and the controller’s increase in cash-buying speed.
In addition to the optimal control described in Subsection 3.3, we evaluate the performance of three reference strategies:
@@
  3. 3.
    The equal weights strategy: Maintains a fixed 50/50 allocation between cash and Treasury Bills, rebalancing at each settlement window to restore the target split and converting bills to cash when immediately necessary to cover redemption requests.
@@ 4.2 Parameter Estimation
Throughout the numerical section we measure time in hours. Annualized rates are converted to hourly continuously-compounded rates by division by $8760$:
  • 90-day Treasury-bill yield: $r_{\mathrm{bill}}=0.054/8760\approx 6.16\times 10^{-6}\ \mathrm{hour}^{-1}.$ (Over an $8$-hour settlement window this corresponds to $r_{\mathrm{bill}}\Delta_{j}\approx 4.93\times 10^{-5}$.)
  • Discount rate: $\rho=0.08/8760\approx 9.13\times 10^{-6}\ \mathrm{hour}^{-1}.$ (Over an $8$-hour window: $\rho\Delta_{j}\approx 7.31\times 10^{-5}$.)
@@
  • Maximum reallocation rate: $\omega_{\max}=0.1\,S_{\mathrm{out}}(0)/\Delta_{j}$, permitting up to 10% of the initial outstanding supply to be shifted per settlement window.
@@
  • Settlement horizon in (6): $\tau_{\mathrm{set}}=\Delta_{j}=8$ hours.
  • Fee bounds (even if $\delta\equiv 0$ in the baseline experiments): $\delta_{\max}=0.01$ (100 bps).
@@
To simplify the interpretation of numerical results, we assume zero mint or redemption fee ($\delta=0$). For modelling the Hawkes processes, we choose parameters to mimic the dynamics observed in major stablecoins (USDC, USDT) during 2024-2025. For baseline intensities, we set $\lambda_{0,R}=80$ events/hour for redemptions and $\lambda_{0,M}=85$ events/hour for mints, corresponding to daily baseline volumes of 1,920 and 2,040 events respectively, which align with typical median volumes for these assets.
@@
To capture the clustering behavior characteristic of stablecoin transactions, we parameterize the self-excitation dynamics with decay rates of $\theta_{R}=2.0$ hour-1 and $\theta_{M}=1.5$ hour-1, yielding half-lives of approximately 20-30 minutes that reflect how quickly bursts of activity dissipate. We set the corresponding jump sizes to $\kappa_{R}=0.8$ and $\kappa_{M}=0.6$, which ensure subcritical branching ratios ($\kappa_{R}/\theta_{R}=0.4<1$ and $\kappa_{M}/\theta_{M}=0.4<1$) necessary for process stability. Finally, transaction sizes follow exponential distributions with mean redemption size $\bar{z}_{R}=\mathdollar 250,000$ and mean mint size $\bar{z}_{M}=\mathdollar 300,000$, values representative of institutional trading volumes in these markets.
The cost function parameters are scaled to be commensurate with hourly dollar cash flows. For peg stability we penalize squared deviations and choose $c_{\mathrm{peg}}$ such that sustaining a deviation of $\Delta P_{\mathrm{tol}}=10$ bps (=0.001=0.001) carries an hourly cost comparable to the hourly T-bill carry on the outstanding supply:
$c_{\mathrm{peg}}=\frac{r_{\mathrm{bill}}\,S_{\mathrm{out}}(0)}{\Delta P_{\mathrm{tol}}^{2}}\approx\frac{(6.16\times 10^{-6})\cdot(10^{10})}{(10^{-3})^{2}}\approx 6.16\times 10^{10}.$ (67)
For fee frictions we choose $c_{\mathrm{fee}}$ such that charging $\delta_{\mathrm{tol}}=10$ bps on a typical hour of redemptions incurs a quadratic penalty comparable to the corresponding fee revenue:
$c_{\mathrm{fee}}=\frac{\lambda_{0,R}\bar{z}_{R}}{\delta_{\mathrm{tol}}}\approx\frac{(100)\cdot(2.5\times 10^{5})}{10^{-3}}\approx 2.5\times 10^{10}.$ (68)
Trading costs reflect treasury-market execution frictions. We set the linear term to a 1bp half-spread, $\lambda_{\omega}=10^{-4}$, so that $\int\lambda_{\omega}|\omega(t)|\,dt$ equals approximately 1bp times traded notional, and $\rho_{\omega}=8\times 10^{-13}$ (in hour/\$ units) for the quadratic term. The peg adjustment speed is set to $\eta=\ln(2)/4\approx 0.173\ \mathrm{hour}^{-1}$, corresponding to a half-life of 4 hours for the gap between $\Delta P(t)$ and $\Delta P_{\mathrm{fair}}(t)$. For the fee effectiveness we set $\gamma=5.0\ \mathrm{hour}^{-1}$. Finally, we set the peg-to-redemption feedback strength $\zeta$ to 2000.
@@
For initial conditions, we consider a system at scale with outstanding supply $S_{\mathrm{out}}(0)=\mathdollar 10$ billion and zero deviation from the peg ($\Delta P(0)=0$). Transaction intensities begin at baseline levels: $\lambda_{R}(0)=\lambda_{0,R}$ and $\lambda_{M}(0)=\lambda_{0,M}$. The reserve allocation strategies differ between controllers: both the optimal controller and the maximum yield controller begin with $R_{\mathrm{liq}}(0)=\mathdollar 1$ billion (10%) in liquid assets and $R_{\mathrm{bill}}(0)=\mathdollar 9$ billion (90%) in Treasury bills, while the maximum liquidity controller starts with all reserves (\$10 billion) allocated to liquid cash.